Carmelita Inc., has the following information available:

Costs from Beginning Inventory Costs from current Period
Direct materials $5,800 $28,000
Conversion costs 6,000 146,600

At the beginning of the period, there were 400 units in process that were 44% complete as to conversion costs and 100% complete as to direct materials costs. During the period, 4,900 units were started and completed. Ending inventory contained 300 units that were 27% complete as to conversion costs and 100% complete as to materials costs. Assume that the company uses the FIFO process cost method. Round cost per unit figures to two cents, i.e., $2.22, when calculating total costs.

The total costs that will be transferred into Finished Goods for units started and completed were

a.$174,460

b.$221,564

c.$120,008

d.$164,395
Answer #1

Equivalent units of production:

Units Material Conversion
% of completion Units % of completion Units
Beginning work in process 400 56% 224
Started and completed 4,900 100% 4,900 100% 4,900
Transferred out 5,000
Ending work in process 300 100% 300 27% 81
Total 5,300 5,200 5,205

Cost per equivalent unit:

Material cost per unit = $28,000 / 5,200 = $5.38

Conversion cost per unit = $146,600 / 5,205 = $28.17

Cost of finished goods transferred for units started and completed :

Direct material = 4,900*$5.38 = $26,362

Conversion cost = 4,900*$28.17 = $138,033

Cost of finished goods transferred for units started and completed = $26,362 + 138,033 = $164,395

Hence option d is correct.